OpenStudy (anonymous):

Help please, Divide (x³-x+8) / (x-2)

OpenStudy (anonymous):

x^3-x+8=x^3-2x^2+2x^2-4x+3x-6+14 =x^2(x-2)+2x(x-2)+3(x-2)+14=(x^2+2x+3)(x-2)+14 (x^3-x+6)/(x-2) =x^2+2x+3+[14/(x-2)]

OpenStudy (anonymous):

...huh?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

im trying to writ this down on paper but... you lost me @ the second step

OpenStudy (anonymous):

see i have rewrite the expression so that i can take (x-2) common.

OpenStudy (anonymous):

see you have to divide with x-2

OpenStudy (anonymous):

i have introduced -2x^2 to make with x^3 as x^2(x-2). and i have add 2x^2 to nullify it...

OpenStudy (anonymous):

i have written -x as -4x+3x. and 8 as -6 +14.

OpenStudy (anonymous):

ghhosst.............

OpenStudy (anonymous):

have you now understood.

OpenStudy (amistre64):

x^2 +2x +3 -------------- x-2 | x^3 -x +8 -x^3 +2x^2 ----------- 2x^2 -x +8 -2x^2 +4x ------------ 3x +8 -3x +6 -------- 14 <- remainder

OpenStudy (anonymous):

yea :) i c

OpenStudy (anonymous):

the way it was written just threw me off, but i see how its done now thank you :)

OpenStudy (anonymous):

how did yu get x^2 +2x +3 ?

OpenStudy (anonymous):

simply we divided

OpenStudy (anonymous):

nvm.. yu put the answer on the top. cool

OpenStudy (anonymous):

see i had x^3-x+8=x^2(x-2)+2x(x-2)+3(x-2)+14=(x^2+2x+3)(x-2)+14 so taking (x-2 common from the first three terms i have (x^2+2x +3)(x-2)